Mar is 4 years older than Ron. Chan is twice as old as Mar. The sum of
their 3 ages is 8 times Ron’s age. How old is Ron now?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

Mar             r+4
Ron             r
Chan            2r+8
SUM             8r

NOT THIS--------------------

Mar             r+3
Ron             r
Chan            2r+6
SUM             8r

misread part of description

Answer by ikleyn(52788)

.
Mar is 4 years older than Ron. Chan is twice as old as Mar. The sum of
their 3 ages is 8 times Ron’s age. How old is Ron now?
~~~~~~~~~~~~~~~~~


            The setup in the post by @josgarithmetic is  INCORRECT.
            I came to solve the problem correctly.


Let x be the Ron's age.

then the Mar's age is (x+4) years;  Chan's age is 2(x+4).


An equation for the total age is

    x + (x+4) + 2(x+4) = 8x.


Simplify and find x

    x + x + 4 + 2x + 8 = 8x

          4 + 8        = 8x - x - x - 2x

            12         = 4x

             x         = 12/4 = 3.


ANSWER.  Ron is 3 years old.

Solved  (correctly).